OpenStudy (anonymous):

FAN AND MEDAL! Ike made the following long jumps - 12 ft, 14.5 ft, 16 ft, and 13.5 ft. What is the mean absolute deviation of his four jumps...

Vocaloid (vocaloid):

first find the mean of those 4 data values

OpenStudy (anonymous):

14

Vocaloid (vocaloid):

excellent, the mean is 14 our next step is to look at each data point and calculate its distance from 14 so basically, we calculate: | 12 - 14 | = __ | 14. 5 - 14 | = __ |16 - 14 | = __ |13.5 - 14| = __

Vocaloid (vocaloid):

| | means absolute value

OpenStudy (anonymous):

-2, .5, 2, -.5

Vocaloid (vocaloid):

almost, absolute value means we take off any negative signs so our values are 2, 0.5, 2, 0.5 now take the mean of those 4 numbers and we'll have our answer!

OpenStudy (anonymous):

5!

OpenStudy (anonymous):

:D

Vocaloid (vocaloid):

not quite, you need to take the mean

Vocaloid (vocaloid):

what is the mean of 2, 0.5, 2, and 0.5?

OpenStudy (anonymous):

1.25

Vocaloid (vocaloid):

yup! that's our answer

OpenStudy (anonymous):

:D thank you

OpenStudy (anonymous):

umm do you know how to interquartile range?